OpenStudy (amy0799):

Find all points on the circle x2 + y2 = 64 where the slope is 3/4. ( , ) (positive y-coordinate) ( , ) (negative y-coordinate)

OpenStudy (amy0799):

typo x^2+y^2=4

OpenStudy (welshfella):

you need to solve the system of equations x^2+y^2=4 and y/x = 3/4

OpenStudy (welshfella):

use the substitution method

OpenStudy (empty):

Are you allowed to use calculus?